# Imagination Augmented Generation: Learning to Imagine Richer Context for Question Answering over Large Language Models

### v3

@@ 3.4 Training with Long Context Distillation
Within the framework of knowledge distillation, components such as hidden representations Jiao et al. (2020), attention dependencies Wang et al. (2020), and relations among representations Park et al. (2021) are regarded as valuable knowledge for transfer. In this paper, we consider long context distillation (LCD) as the contextualized knowledge that mainly guides the student.
Specifically, the teacher model FiD Izacard and Grave (2021), which utilizes longer contextual inputs and theoretically contains more information (richer context). It will activate more specific internal knowledge and serve as a supervisory model. The teacher model assists the student model T5 Roberts et al. (2020a), which has the same size as the teacher and leverages short contextual inputs. This aids in activating richer feature representations and knowledge. The optimization objective for the student model at each mini-batch $z_{r}=(x_{r},y_{r})$ is:
@@
As shown in the right of Figure 2, we perform an additional representation alignment for better knowledge transfer. In our distillation, both teacher model and student model have the $L$ layers, we feed the text into them and can obtain the corresponding output hidden states $\{H_{l}^{t}\}_{l=0}^{L}$, $\{H_{l}^{s}\}_{l=0}^{L}$, and attention matrices $\{A_{l}^{t}\}_{l=1}^{L}$, $\{A_{l}^{s}\}_{l=1}^{L}$. We suppose the student’s $l$-th layer is aligned with the teacher’s $l$-th layer, then the outputs of the student (i.e., $H_{l}^{s}$ and $A_{l}^{s}$) should be close to the teacher’s (i.e., $H_{l}^{t}$ and $A_{l}^{t}$). For aligning hidden states, following Park et al. (2021), we use cosine distance COS to calculate the proximity between the hidden states of the teacher and the student:
@@
The overall objective for training IMcQA is the weighted sum of the two objectives:
@@
IMcQA (Ours) 10 49.9 50.9 69.7 70.3 51.5 52.8
IMcQA (Ours) 30 53.1 - 70.5 - 52.0 -
@@
IMcQA (Ours) 10 48.8 49.2 70.9 72.2 54.5 55.6
# Imagination Augmented Setting
@@ 4 Experiment
In this section, we conduct experiments to demonstrate the effectiveness and efficiency of IAG and IMcQA on QA. The experiment mainly answers four research questions (RQs):
RQ1: Can IAG and IMcQA achieve knowledge augmentation for QA over LLMs? (§ 4.4)
RQ2: Does IMcQA have a good out-of-distribution generalization ability? (§ 4.5)
RQ3: Does IAG have advantages in effectiveness and efficiency compared to RAG and GAG? (§ 4.6)
RQ4: What is the role of explicit and implicit imagination modules in IMcQA? (§ 4.7)
@@ 4.2 Baselines
[…]
To demonstrate the plug-and-play capability of IMcQA on the zero-shot settings of LLMs ($\geq$ 3B), we use Llama2-7B and -13B Touvron et al. (2023) as the basic model. We evaluate with 6 diverse settings: without retrieval, with retrieval, with LoRA, RECITE Sun et al. (2023), HICL Wang et al. (2024) and using the proposed IMcQA.
@@ 4.3 Implementations
In the pretraining stage, the context generator initialized with T5-large utilizes the generated question-compressed pairs. During the second stage, the teacher model employs a FiD reader with different sizes (FiD-l and FiD-xl) that are fine-tuned on the training split of target datasets. The student model freezes the backbone and updates solely the hypernetwork [^5] B.3 contains more implementation and baseline details..
@@ 4.4.1 Supervised Setting
Table 1 shows the performance results, full results including T5-Base are in the C.1. When juxtaposed with closed-book models, RAG, and GAG methods, our proposed IAG framework IMcQA method exhibits state-of-the-art (SOTA) performance with the equivalent magnitude of document numbers.
In the closed-book setting (in the upper part of the table), our method outperforms the baseline by an average of +2% EM score, indicating its excellence in utilizing internal knowledge with imagination. It’s especially noteworthy that as the model size expands, the performance advantages of the imagination become ever more evident.
The following three parts show the experimental results under the open domain setting [^6] Due to memory constraints from online distillation, we only compare the IMcQA-l under the RAG setting using 30 documents with FiD using 100 documents.. Although our method only deals with one short dummy document, it can still achieve results similar to or better than the RAG and GAG methods, which handle 10 documents. The findings reveal that IMcQA exploits imagined condensed text to strike a balance between efficiency and overhead. Beyond this, when IMcQA implements 10 retrieved documents under RAG, it overtakes RFiD’s performance, posting 1.6% higher in NQ, 4.4% in TQA, and 2.7% in WQ. IMcQA employing 10 generated documents under GAG outstrips GENREAD (clustering)’s performance, documenting a 4.5% increment in NQ, 0.7% in TQA, and 1.1% in WQ.
Figure 4: Zero-Shot results (EM, %) of Llama2-7B on three open-domain QA datasets. The number in parentheses indicates the number of documents used.
@@ 4.4.2 Zero-shot Setting
Figure 4 illustrates the zero-shot results for LLMs implementing IMcQA with a frozen Llama2-7B and -13B. This research seeks to explore the possibility of enhancing LLMs via IAG. Due to the high computational demands of training, we only fine-tuned the hypernetwork on a mixed dataset without LCD in this experiment and evaluated performance in a zero-shot setting [^7] Detailed prompts information can be found in the B.2.. More zero-shot setting results can be seen in C.3.
We discerned that Llama2’s performance can be enhanced by imagining knowledge autonomously. While leveraging explicit imagined context could amplify the average EM +1%, this is not as significant as the improvement achieved by retrieving 10 documents, indicating the limitations of relying solely on prompt cues for triggering corresponding knowledge. IMcQA can enhance knowledge via two main imagination processes, escalating EM by +15.33% for NQ, +11.97% for TQA, and +16.38% for WQ. Compared to two other advanced RAG methods, IMcQA using a single document, performs only 1 EM lower than the HICL method Wang et al. (2024) on the TQA but achieves +10% EM on the NQ and +5% EM on the WQ. With IAG, Llama2-7B demonstrated an average improvement of +14% across the three datasets. This trend is also observed in Llama2-13B’s results (Figure 6). This implies that even in zero-shot settings, our method can still offer substantial benefits to LLMs.
@@
IMcQA 10 47.01 55.74 24.13 49.92 60.03 25.79
LoRA 1 5.43 3.15 4.02 4.42 6.50 7.87
IMcQA 1 40.14 46.61 18.92 42.32 54.80 22.05
Table 2: IID and OOD results. The performance on three open-domain datasets for the model trained on NQ is reported, with the underlined values indicating the IID performance. Full OOD results and details of three datasets are shown in the C.2.
@@ 4.5 Out-Of-Distribution (OOD) Performance
To further demonstrate the generalizations of the IMcQA method and the importance of hypernetwork, we also evaluate its performance in OOD generalizations. Table 2 shows the IID and OOD performance of FiD, and IMcQA methods with different document settings when training on NQ (From NQ generalization to the other two datasets).
It is patently clear that an increment in document provision leads to better OOD performance, likely due to the presence of answer-oriented content within these documents. Remarkably, IMcQA can come within a relatively narrow 5% gap of FiD, even when utilizing a single imagined document as opposed to 10 retrieved documents.
Simultaneously, IMcQA generally showcases superior performance in OOD when provided with 10 retrieved documents. This superiority can be traced back to the pivotal role played by hypernetwork in generating LoRA adapters’ weights based on questions. This equips models with the capability to invoke and access internal knowledge based on context-specific discourse rather than confining to resolving distinct questions.
Models Training Params # Documents # Avg Tokens Inference Time GPU Memory Training Time
T5 220M 0 19.8 79.8s 2828M 0.9h
IMcQA 139.3M 0 19.8 82.3s 2710M 1.2h
IMcQA 139.3M 1 522.1 214.6s 2882M 1.7h
FiD 220M 10 1748.3 683.3s 4358M 2.3h
GENREAD 220M 10 1912.5 704.8s 4412M -
FiD 220M 100 16625.7 1293.2s 19048M 5.8h
Table 3: Training and inference cost on the NQ. The backbone model is T5-Base.
@@ 4.6 Training Cost and Inference Speed-up
We proceeded to measure the inference speed documented in GPU time and training time for 5000 steps on the NQ dataset using T5-Base. The experiments were conducted on a single RTX 3090 GPU, maintaining a standard batch size of 8 during training and 1 during inference.
@@
As evident from Table 3, the proposed method’s advantage lies in its diminished requirement for parameter updates, which can be attributed to the shared hypernetwork’s utilization that generates LoRA adapters, thereby negating the necessity of individual LoRA adapters’ setup. Despite the lack of a training advantage due to distillation constraints, IMcQA achieves efficient reasoning through an extremely lightweight design, saving more than half the training time compared to methods using a large number of documents (×$\times 0.3$). Compared to the other two methods, the processed tokens are significantly decreased, while either outperforming them or showing negligible differences in performance. This represents an optimal trade-off between efficiency and computational demand. Moreover, unlike GAG, our approach incurs no financial costs associated with API calls, and the reduced model size facilitates faster generation.
4.7 Ablation Experiment
@@
This study introduces two key imagination processes to stimulate LLMs’ internal knowledge: Explicit Imagination (EI) and Implicit Imagination (II). We particularly examined the influence of different imagination types on performance.
@@
Figure 5 demonstrates that both EI and II are important for IMcQA. Omitting either one results in a considerable reduction in performance, with a drop exceeding 10% observed when EI is neglected. This is harmonious with the initial observation that performance improvement becomes more noticeable when relevant documents are available, thus underscoring EI’s superiority.
@@
The outcomes of Long Context Distillation (LCD) and the application of EI in the hypernetwork also make marginal contributions to the overall results. This validates the previous assertion that a more extensive context tends to optimize performance, although with limited gains. The experiments and analysis above demonstrate the importance of each component and the effectiveness of our IMcQA method.
Figure 5: Ablation experiment results (%) of T5-Base on three open-domain QA datasets.
5 Conclusion and Future Work
@@
This study proposes a novel knowledge-augmented strategy for LLMs, namely Imagine Augmented Generation (IAG), and a specific method IMcQA for open domain question answering. The proposed method effectively activates and utilizes intrinsic knowledge within LLMs through two imaginations: explicit imagination, and implicit imagination. Experimental results demonstrate a significant improvement in QA performance while remaining relatively lightweight. Although the main focus of this method is on one specific task, we believe these findings can offer a novel perspective on how to better harness the potential of LLMs. In the future, we plan to apply IAG to more NLP tasks and explore multimodal knowledge-augmented generation.
@@
Method.   Our method relies on the knowledge learned by LLMs in the pre-training phase, which may limit the model’s ability to quickly adapt to new information. The dependency on internal knowledge activation in IAG may lead to a less transparent decision-making process in the model, making it challenging to explain the logic behind the generated answers. In the future, there is a need to continue exploring adaptive knowledge enhancement methods to optimize results further.
@@ A.1 Comparison of Three Paradigms
Next, we compare IAG, RAG, and GAG across four criteria for a more intuitive understanding. From the table 4, it can be observed that the document relevance obtained by IAG and GAG is higher, while RAG heavily relies on the retriever and external knowledge base. In terms of document length usage, IAG only needs to use a virtual document, greatly reducing the number of tokens. Therefore, IAG is superior to the other two methods in terms of reasoning time.
@@ A.2 Context Generator
There are two main goals in the pre-training of the model in the first stage of IAG (context generator): first, to improve its ability as a document generator by learning to generate rich and concise documents; second, to introduce some external knowledge that the model initially does not possess. It’s worth noting that the second goal is crucial, as the model may encounter knowledge it has not yet learned. Thus, IAG does not rely on external large models or retrievers for external reasoning and can complete reasoning independently.

### v4

@@ 3.4 Training with Long Context Distillation
Within the knowledge distillation framework, elements like hidden representations Jiao et al. (2020), attention dependencies Wang et al. (2020), and relationships among representations Park et al. (2021) are considered essential for effective knowledge transfer. In this paper, we introduce long context distillation (LCD) as the contextualized knowledge that primarily guides the student model. Specifically, the teacher model, FiD Izacard and Grave (2021), which processes longer contextual inputs, theoretically contains more information due to its richer context. This enables it to activate more specific internal knowledge, serving as a supervisory model. The teacher model aids the student model, T5 Roberts et al. (2020a), which is of the same size but uses shorter contextual inputs, in activating richer feature representations and knowledge. The optimization objective for the student model at each mini-batch $z_{r}=(x_{r},y_{r})$ is:
@@
As illustrated on the right of Figure 2, we perform additional representation alignment to facilitate better knowledge transfer. In our distillation process, both the teacher and student models have $L$ layers. The input text is processed through these layers, yielding corresponding output hidden states $\{H_{l}^{t}\}_{l=0}^{L}$ and $\{H_{l}^{s}\}_{l=0}^{L}$, along with attention matrices $\{A_{l}^{t}\}_{l=1}^{L}$ and $\{A_{l}^{s}\}_{l=1}^{L}$. For aligning hidden states, we calculate the proximity between the teacher’s and student’s hidden states using cosine distance (COS) Park et al. (2021).
@@
The overall objective for training AAG is the weighted sum of the two objectives:
@@
AAG (Ours) 10 49.9 50.9 69.7 70.3 51.5 52.8
AAG (Ours) 30 53.1 - 70.5 - 52.0 -
@@
AAG (Ours) 10 48.8 49.2 70.9 72.2 54.5 55.6
# Awakening Augmented Setting
@@ 4 Experiment
In this section, we conduct experiments to demonstrate the effectiveness and efficiency of AAG on QA. The experiment mainly answers four research questions (RQs):
RQ1: Can AAG achieve knowledge augmentation for QA over LLMs? (§ 4.4)
RQ2: Does AAG have a good out-of-distribution generalization ability? (§ 4.5)
RQ3: Does AAG have advantages in effectiveness and efficiency compared to RAG and GAG? (§ 4.6)
RQ4: What is the role of explicit and implicit awakening modules in AAG? (§ 4.7)
@@ 4.2 Baselines
Both the moderately sized language model (<1B) and the large language model ($\geq$ 3B) are under consideration. T5 Roberts et al. (2020a) is selected as the backbone for our moderately sized language models. We evaluate our proposed AAG against several knowledge-enhanced approaches, which include RAG models such as DPR Karpukhin et al. (2020), RAG Lewis et al. (2020), EAR Chuang et al. (2023), RFiD Wang et al. (2023a), FILCO Wang et al. (2023d) and FiD Izacard and Grave (2021), as well as the GAG model GENREAD Yu et al. (2023), and parameters efficient fine-tuning method LoRA Hu et al. (2021).
To demonstrate the plug-and-play capability of AAG on the zero-shot settings of LLMs ($\geq$ 3B), we use Llama2-7B and -13B Touvron et al. (2023) as the basic model. We evaluate with 6 diverse settings: without retrieval, with retrieval, with LoRA, RECITE Sun et al. (2023), HICL Wang et al. (2024) and using the proposed AAG.
@@ 4.3 Implementations
In the pretraining stage, the context generator initialized with T5-large utilizes the generated question-compressed pairs. During the second stage, the teacher model employs a FiD reader with different sizes (FiD-l and FiD-xl) that are fine-tuned on the training split of target datasets. The student model freezes the backbone and updates solely the hypernetwork, FFN and norm layers. B.3 contains more implementation and baseline details.
@@ 4.4.1 Supervised Setting
Table 1 presents the performance results, with full results including T5-Base detailed in C.1. Compared to closed-book models, as well as RAG and GAG methods, our proposed AAG method, achieves state-of-the-art (SOTA) performance using an equivalent number of documents.
In the closed-book setting (upper part of the table), our method surpasses the baseline by an average of +2% EM score, demonstrating its superior ability to leverage internal knowledge through awakening. Notably, as the model size increases, the performance gains from the awakening approach become even more pronounced.
The following sections present the experimental results in the open domain setting [^4] Due to memory constraints, AAG under the RAG setting using 30 documents.. Notably, proposed AAG using just one short dummy document, matches or exceeds the performance of RAG and GAG methods, which process 10 documents. These results demonstrate that AAG effectively balances efficiency and overhead by leveraging imagined compressed text.
AAG outperforms baselines when documents-matched. When AAG utilizes 10 retrieved documents under RAG setting, it surpasses RFiD performance by 1.6% in NQ, 4.4% in TQA, and 2.7% in WQ. When AAG utilizes 10 generated documents under the GAG setting, it surpasses strong baseline GENREAD (clustering) performance by 4.5% in NQ, 0.7% in TQA, and 1.1% in WQ.
Figure 4: Zero-Shot results (EM, %) of Llama2-7B on three open-domain QA datasets. The number in parentheses indicates the number of documents used. More zero-shot setting results can be seen in C.3.
@@ 4.4.2 Zero-shot Setting
Figure 4 illustrates the zero-shot results for LLMs implementing AAG with a frozen Llama2-7B and -13B. This research seeks to explore the possibility of enhancing LLMs via AAG. Due to the high computational demands of training, we only fine-tuned the hypernetwork on a mixed dataset without LCD in this experiment and evaluated performance in a zero-shot setting. Detailed prompt information can be found in the B.2.
We discerned that Llama2’s performance can be enhanced by imagining knowledge autonomously. While leveraging explicit imagined context could amplify the average EM +1%, this is not as significant as the improvement achieved by retrieving 10 documents, indicating the limitations of relying solely on prompt cues for triggering corresponding knowledge. AAG can enhance knowledge via two main awakening processes, escalating EM by +15.33% for NQ, +11.97% for TQA, and +16.38% for WQ. Compared to two other advanced RAG methods, AAG using a single document performs only 1 EM lower than the HICL method Wang et al. (2024) on the TQA but achieves +10% EM on the NQ and +5% EM on the WQ. With AAG, Llama2-7B demonstrated an average improvement of +14% across the three datasets. This trend is also observed in Llama2-13B’s results (Figure 5). This implies that even in zero-shot settings, our method can still offer substantial benefits to LLMs.
@@
AAG 10 47.01 55.74 24.13 49.92 60.03 25.79
LoRA 1 37.17 45.20 15.62 37.61 48.50 20.71
AAG 1 40.14 46.61 18.92 42.32 54.80 22.05
Table 2: IID and OOD results. The performance on three open-domain datasets for the model trained on NQ is reported, with underlined values indicating IID performance. Full OOD results and details of the three datasets are provided in the C.2.
@@ 4.5 Out-Of-Distribution (OOD) Performance
To further demonstrate the generalizations of the AAG method and the importance of hypernetwork, we also evaluate its performance in OOD generalizations. Table 2 shows the IID and OOD performance of FiD, and AAG methods with different document settings when training on NQ (From NQ generalization to the other two datasets).
It is patently clear that an increment in document provision leads to better OOD performance, likely due to the presence of answer-oriented content within these documents. Remarkably, AAG can come within a relatively narrow 5% gap of FiD, even when utilizing a single imagined document as opposed to 10 retrieved documents.
Simultaneously, AAG generally showcases superior performance in OOD when provided with 10 retrieved documents. This superiority can be traced back to the pivotal role played by hypernetwork in generating LoRA adapters’ weights based on questions. This equips models with the capability to invoke and access internal knowledge based on context-specific discourse rather than confining to resolving distinct questions.
Models Training Params # Docs # Avg Tokens Inference Time Training Time
T5 220M 0 19.8 79.8s 0.9h
AAG 139.3M 0 19.8 82.3s 1.2h
AAG 139.3M 1 522.1 214.6s 1.7h
FiD 220M 10 1748.3 683.3s 2.3h
GEN. 220M 10 1912.5 704.8s -
FiD 220M 100 16625.7 1293.2s 5.8h
Table 3: Training and inference cost on the NQ.
@@ 4.6 Training Cost and Inference Speed-up
We proceeded to measure the inference speed documented in GPU time and training time for 5000 steps on the NQ dataset using T5-Base. The experiments were conducted on a single RTX 3090 GPU, maintaining a standard batch size of 8 during training and 1 during inference. A detailed inference case is shown in the Appendix D.
@@
As evident from Table 3, the proposed method’s advantage lies in its diminished requirement for parameter updates, which can be attributed to the shared hypernetwork’s utilization that generates LoRA adapters, thereby negating the necessity of individual LoRA adapters’ setup. Despite the lack of a training advantage due to distillation constraints, AAG achieves efficient reasoning through an extremely lightweight design, saving more than half the training time compared to methods using a large number of documents (0.3×0.3$\times$0.3 ×). Compared to the other two methods, the processed tokens are significantly decreased, while either outperforming them or showing negligible differences in performance. This represents an optimal trade-off between efficiency and computational demand. Moreover, unlike GAG, our approach incurs no financial costs associated with API calls, and the reduced model size facilitates faster generation.
4.7 Ablation Study
@@
This study introduces two key awakening processes to stimulate LLMs’ internal knowledge: explicit awakening (EA) and implicit awakening (IA). We particularly examined the influence of different awakening types on performance.
@@
Table 4 demonstrates that both EA and IA are important for AAG. Omitting either one results in a considerable reduction in performance, with a drop exceeding 30% observed when EA is neglected. This is harmonious with the initial observation that performance improvement becomes more noticeable when relevant documents are available, thus underscoring EA’s superiority.
@@
The outcomes of Long Context Distillation (LCD) including $\mathcal{L}_{s}$ and $\mathcal{L}_{align}$ also make marginal contributions to the overall results. This validates the previous assertion that a more extensive context tends to optimize performance, although with limited gains. The impact of EA on the application of hypernetworks is minimal (<3%), indicating that hypernetworks in IA primarily serve to awaken parameter knowledge rather than to utilize the generated context. The experiments and analysis above demonstrate the importance of each component and the effectiveness of our AAG method.
Methods # Docs ($\downarrow$) NQ ($\uparrow$) TQA ($\uparrow$) WQ ($\uparrow$)
AAG 1 40.14 60.75 41.73
w/o EA 0 23.89 ($\downarrow 40\%$) 22.69 ($\downarrow 63\%$) 30.31 ($\downarrow 27\%$)
In. w/o EA 1 38.85 ($\downarrow 3\%$) 59.62 ($\downarrow 2\%$) 40.65 ($\downarrow 3\%$)
w/o IA 1 33.48 ($\downarrow 17\%$) 51.19 ($\downarrow 16\%$) 34.72 ($\downarrow 17\%$)
w/o LCD 1 33.96 ($\downarrow 15\%$) 53.27 ($\downarrow 12\%$) 29.39 ($\downarrow 29\%$)
w/o $\mathcal{L}_{s}$ 1 34.24 ($\downarrow 14\%$) 54.90 ($\downarrow 10\%$) 31.67 ($\downarrow 24\%$)
w/o $\mathcal{L}_{align}$ 1 37.41 ($\downarrow 7\%$) 56.38 ($\downarrow 7\%$) 39.26 ($\downarrow 6\%$)
Table 4: Ablation studies on three open-domain QA datasets. The backbone model is the T5-base. "In." means the input of the hypernetwork § 3.3.
5 Conclusion
@@
This study proposes a novel knowledge-augmented strategy for Large Language Models (LLMs), namely Awakening Augmented Generation (AAG) for open domain question answering. The AAG effectively harnesses the inherent knowledge of the LLMs through a dual-awakening approach to awaken a richer context. Explicit awakening with the context generator generates a short dummy document as symbolic context, while implicit awakening uses hypernetwork to convert the question and the document into adapters inserted into the LLMs as parameter context. Experimental results demonstrate a significant improvement in performance while remaining relatively lightweight. Although the main focus of this method is on one specific task, we believe these findings can offer a novel perspective on how to better harness the potential of LLMs.
@@
Method.   Our method relies on the knowledge learned by LLMs in the pre-training phase, which may limit the model’s ability to quickly adapt to new information. The dependency on internal knowledge activation in AAG may lead to a less transparent decision-making process in the model, making it challenging to explain the logic behind the generated answers. In the future, there is a need to continue exploring adaptive knowledge enhancement methods to optimize results further.
Hypernetwork.   For lightweight and efficient settings, our hypernetwork employs a two-layer MLP. However, some studies use larger models, such as GPT-2 or T5, as hypernetworks. Due to computational resource constraints, we did not explore or compare the effects of different hypernetwork models on the results. Nonetheless, our method primarily focuses on generating parameter-efficient modules to enhance knowledge activation and generalization.
@@ A.1 Comparison of Three Paradigms
Next, we compare AAG, RAG, and GAG across four criteria for a more intuitive understanding. From the table 5, it can be observed that the document relevance obtained by AAG and GAG is higher, while RAG heavily relies on the retriever and external knowledge base. In terms of document length usage, AAG only needs to use a virtual document, greatly reducing the number of tokens. Therefore, AAG is superior to the other two methods in terms of reasoning time.
@@ A.2 Context Generator
There are two main goals in the pre-training of the model in the first stage of AAG (context generator): first, to improve its ability as a document generator by learning to generate rich and concise documents; second, to introduce some external knowledge that the model initially does not possess. It’s worth noting that the second goal is crucial, as the model may encounter knowledge it has not yet learned. Thus, AAG does not rely on external large models or retrievers for external reasoning and can complete reasoning independently.